OpenStudy (anonymous):

solve for x e^2x-e^x+(1/4)=0

OpenStudy (anonymous):

like solving \[z^2-z+\frac{1}{4}=0\]

OpenStudy (anonymous):

then put \[z=e^x\] and take the log to get your answer. do you need the details?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

as this is a perfect square you get \[(z-\frac{1}{2})^2=0\] so \[z=\frac{1}{2}\] making \[e^x=\frac{1}{2}\]and thus \[x=\ln(\frac{1}{2})\]

OpenStudy (anonymous):

hope the steps are clear. if you want (not necessary) you can write \[\ln(\frac{1}{2})=-\ln(2)\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

can you help me with another question?